OpenStudy (anonymous):

solve 12^(x-2)=20 x=____?

OpenStudy (anonymous):

240

OpenStudy (anonymous):

apply ln on both side & then use this rule ln(a)^b = b * ln(a)

OpenStudy (anonymous):

can you plug numbers into equation then i will claculate, this is a foreign language to me

OpenStudy (anonymous):

would a=12?

OpenStudy (anonymous):

Yes a=12 & b = (x - 2)

OpenStudy (anonymous):

ln(12)^x-2 = x-2 * ln(12)

OpenStudy (anonymous):

yes correct so we have (x-2)*ln(12) =ln(20) right ?

OpenStudy (anonymous):

i need answer in decimal

OpenStudy (anonymous):

first find move ln(12) to right hand side & then plug values of ln(12) & ln(20) 7 simplify it